# Efficiency Meets Fidelity: A Novel Quantization Framework for Stable Diffusion

### v1

@@ 4.2 Components For Higher Fidelity
Low memory time information precalculation. In a Stable Diffusion model, the time-step $t$ is firstly encoded by time-embedding layers, then passed through time-projection layers in each Bottleneck block. The Time information $e_{p}$ inserted into the UNet is calculated as follows:
$e_{t}=emb(t),ep_{t,i}=proj_{i}(e_{t})$ (8)
@@
We observe that the quantization of both the time embedding and time embedding projection layers of the model has a significant impact on the quality of the generated images. When the inference configuration is determined, the output of the time embedding module $e_{t}$ is only related to timesteps, and $ep$ is dependent only on $e_{t}$. Consequently, $ep$ is finite and invariant. Therefore, we remove time-embedding and time-projection from the model and save the $ep$, which is directly input into the Resnet blocks of the model. The memory usage and computational cost of $ep$ are much smaller than the parameters of the time embedding module and the time embedding projection layers.
Mixed-Precision Quantization Strategy. Recent studies  [57, 20] observe that compressing different blocks in Diffusion models can lead to different image generation quality. However, quantizing a certain block, fine-tuning, and then evaluating the model, are computationally intensive and time-consuming. Moreover, the sensitivity of each block is coarse-grained, which may lead to suboptimal compression. To address these issues, we propose a mixed-precision quantization strategy to identify sensitive layers and assign different bit-width to different layers based on their sensitivity.
Our objective is to evaluate the sensitivity $S_{i}$ of each layer. We consider i-th layer (e.g. linear layer) of the model with weight ${W_{i}}\in\mathbf{R}^{C_{\text{in}}\times C_{\text{out}}}$, given dataset $\mathcal{D}=\left\{x_{m},y_{m}\right\}_{m=1}^{n}$ comprising $\mathbf{n}$ samples.Specifically, to estimate $S_{i}$, the deviation in the loss function caused by $W_{i}$ from original value to zero can be formulated as :
$S_{i}=\left|\frac{\partial\mathcal{L}\left(\mathcal{D}\right)}{\partial W_{i}}^{\top}W_{i}+\frac{1}{2}W_{i}^{\top}HW_{i}+\mathcal{O}\left(\left\|W_{i}\right\|^{3}\right)\right|$ (9)
where $H$ is the Hessian matrix and $\mathcal{L}$ denotes the loss function. However, the formula cannot be directly computed, since the computation of $H$ on the model is impractical. By employing the Fisher information matrix approximation [3], the computation of loss deviation can be rewritten as:
$S_{i}\approx\left|\frac{\partial\mathcal{L}\left(\mathcal{D}\right)}{\partial W_{i}}^{\top}W_{i}-\frac{1}{2}\left(\frac{1}{n}\sum_{m=1}^{n}\frac{\partial\mathcal{L}(\mathcal{D}_{n})}{\partial W_{i}}W_{i}\right)^{2}\right|$ (10)
where the redundant term can be neglected. Finally, the mix-precision quantization is then conducted based on the sensitivity ranking. We selected the top 5% of layers with the highest sensitivity as sensitive layers, while the bottom 5% of layers are designated as insensitive layers. For A8 quantization, sensitive layers are set to A16, while insensitive layers are set to A4. For W4 quantization, sensitive layers are set to W8. The detailed average bit-width will be displayed in the experiment section.
4.3 Objective Function
@@
We optimize quantized UNet $\epsilon_{q}$ to mimic the output of the floating-point UNet $\epsilon_{fp}$. Given the latent $\{x_{t1}\ldots x_{tn}\}$ at timestep $\{t_{1}\ldots t_{n}\}$, text embedding $\{p_{1}\ldots p_{n}\}$ from frozen text encoder, and precalculated projection $ep$, the output loss is defined as the mean squared error between the quantized and floating-point UNet outputs:
$\mathcal{L}_{\text{out}}=\mathbb{E}\left[\left\|\epsilon_{fp}(x_{t_{1}\ldots t_{n}},p_{1\ldots n},e_{p})-\epsilon_{q}(x_{t_{1}\ldots t_{n}},p_{1\ldots n},e_{p})\right\|_{2}^{2}\right]$ (11)
where $\epsilon_{fp}$ and $\epsilon_{q}$ indicate the floating-point UNet and the quantized UNet, respectively.
@@
The loss function of the feature maps by the sensitive layers is added to ensure they receive more attention:
$\mathcal{L}_{\text{sen}}=\mathbb{E}\left[\left\|f_{fp}^{s}(x_{t_{1}\ldots t_{n}},p_{1\ldots n},e_{p})-f_{q}^{s}(x_{t_{1}\ldots t_{n}},p_{1\ldots n},e_{p})\right\|_{2}^{2}\right]$ (12)
@@
5 Experiments
5.1 Experimental Setup
@@
Datasets. In this paper, we conduct experiments using two distinct datasets: COCO [26] and Stable-Diffusion-Prompts. We utilize prompts from the COCO training dataset to construct the latent dataset. In terms of evaluation, the process is twofold. Following [48], firstly 5,000 prompts are selected from the COCO validation dataset, which has been extensively employed in previous studies. Secondly, an additional 5,000 prompts from the Stable-Diffusion-Prompts dataset are used to assess the generalization capabilities of our quantized model in different prompt scenarios.
@@
Metrics. We evaluate the generative results of the quantized model from the perspectives of distributional similarity and visual similarity. For distributional similarity, we refer to the FID-to-FP [48], which is the Fréchet Inception Distance between images generated by the quantized model and floating-point model. For visual similarity, we consider the commonly used metrics of SSIM [52], LPIPS [62], and PSNR [10]. Additionally, we use CLIP score [13] to evaluate the matching degree between images and prompts. The evaluation code is adopted from  [49] and  [4].
@@
Baselines and implementation. We compare our proposed approach against advanced techniques: Q-diffusion [24], PTQ4DM [44], and PCR [48]. Results are obtained from  [48] or reproduced. We employ the Stable Diffusion v1-4 (resolution of 512x512), Stable Diffusion v2-1 (resolution of 512x512), and the Stable Diffusion XL 1.0 (resolution of 768x768), both sourced from Hugging Face. We compare exclusively with PCR [48] and align the quantization and generation settings with it. Except for special declaration, the standard setup involves a 50-step PNDM sampling process for the Stable Diffusion model and a 50-step Euler sampling process for the Stable Diffusion XL model, with both configurations using a Classifier-Free Guidance (CFG) scale of 7.5. All experiments are conducted using a single NVIDIA A100.
5.2 Dataset Generation Analysis
@@
First of all, we discussed the trade-off of some crucial hyperparameters in the latent dataset generation.
@@
50steps/prompt 4000 6G 1.8h 10.12
1steps/prompt 20000 0.6G 4h 10.14
Table 2: Comparison of different sampling strategies for dataset generation.
More prompts or more timesteps?In the dataset generation process, we can randomly sample varying amounts of latent for each prompt. For comparison, two datasets have been constructed. The first dataset comprises 4000 prompts, with 50 latent per prompt. The second consists of 20000 prompts, each prompt with just 1 latent. As demonstrated in Tab. 2, despite the first dataset having 10$\times$ more latent, it exhibits a similar FID-to-FP. We can infer that a dataset with more prompts is more resistant to overfitting and has a smaller size but with a longer generation time. Given the comparable outcomes of the two strategies, users can select their sampling strategies based on their time requirements or storage requirements.
Figure 5: Comparison of Loss and FID-to-FP Curves under Different Prompt Quantities.
@@
Ours Gen Data (Once for all) (a) (b) (a) (b)
3h $\approx$7.5h 2.5h $\approx$5.5h
@@
Table 3: Efficiency comparison on Stable Diffusion v1-4 and Stable Diffusion XL. (a) denotes sampling 50 steps per prompt, and (b) denotes sampling 1 step per prompt.

### v2

@@ 4.2. Components For Higher Fidelity
Time Information Precalculation. In a Stable Diffusion model, the time-step $t$ is firstly encoded by time-embedding layers, then passed through time-projection layers in each Bottleneck block. The Time information $e_{p}$ inserted into the UNet is calculated as follows:
(8) $e_{t}=emb(t),ep_{t,i}=proj_{i}(e_{t})$
@@
Consistent with prior work (Huang et al., 2024), we observe that quantizing time embedding and time projection layers downgrades the quality of the generated images. While TFMQ (Huang et al., 2024) choose to separately optimize these layers, we propose a more fundamental solution. When the inference configuration is determined, the output of the time embedding module $e_{t}$ is only related to timesteps, and $ep$ is dependent only on $e_{t}$. Consequently, $ep$ is finite and invariant. Therefore, we remove time-embedding and time-projection from the U-Net and save the $ep$, which is directly input into the model, as demonstrated in Fig. 5 (B). The memory usage and computational cost of $ep$ are much smaller than the parameters of the time embedding module and the time embedding projection layers.
Inter-layer Distillation. To enhance generation consistency, we incorporate inter-layer distillation with careful consideration of architectural sensitivities observed in prior work (Wang et al., 2024a; Sui et al., 2024; Liu et al., 2024). For computational efficiency while preserving performance, we selectively apply distillation to some critical layers in each architecture: (1) in U-Net-based models (SD1.4/2.1/XL), we target shortcut connections and projection layers in feed-forward networks; (2) for MMDiT-based models (SD3), we focus on projection layers in the feed-forward blocks of both flow.
Selective Freezing. We analyze oscillations (Nagel et al., 2022) of quantized weight in layers chosen for feature distillation and observe a significant increase in instability after inter-layer distillation is applied, as shown in Fig. 6. To maintain training stability while minimizing computational overhead, we implement a selective iterative freezing mechanism with two key adaptations: (1) reduced freezing frequency (every 500 iterations instead of per-iteration as in (Nagel et al., 2022)), and (2) targeted application to only the mapping layers. This optimized approach introduces negligible runtime overhead (¡5% in SD models) while effectively suppressing distillation-induced oscillations.
Figure 6. Weight oscillations in selected sensitive layers. The oscillation rates significantly arise after adopting inter-layer distillation, while reduced by iterative freezing.
Table 2. Quantitative quality comparisons across different models. Our results outperform other approaches and NF4 baselines. $\dagger$ means reproduced results on our machine. * denotes results in the format of (PCR/Ours). $\downarrow$ means lower is better. $\star$ denotes the key metric. IR is calculated over the first 1000 generated images, while other metrics are calculated over 5000 generated images.
Model Reso. Precision Methods COCO Prompt SD prompt
Similarity Quality Matching Similarity Quality Matching
[…]
SD14 512 FP16 - 0 1 0.16 26.48 0 1 0.17 28.79*/27.28*
W8A8 Q-Diffusion 18.64 - - 26.15 16.22 - - -1.60
W8A8 PTQ4DM 14.60 - - 26,33 13.25 - - -1.02
W8A8.4 PCR 8.35 0.65$\dagger$ 0.12$\dagger$ 26.47 9.52 0.67$\dagger$ 0.12$\dagger$ -0.05
W8A8 Ours 8.57 0.69 0.15 26.55 10.19 0.70 0.14 -0.14
W4A8 Q-Diffusion 20.42 - - 26.15 17.43 - - -1.48
W4A8 PTQ4DM 17.73 - - 26.25 17.28 - - -1.40
W4A8.4 PCR 14.2 0.47$\dagger$ -0.05$\dagger$ 26.48 17.9 0.50$\dagger$ -0.28$\dagger$ -0.74
W4A8 Ours 9.46 0.63 0.16 26.49 12.2 0.62 0.11 -0.09
SD21 512 FP16 - 0 1 0.32 26.07 0 1 0.14 25.48
W4A8.4 PCR 35.9$\dagger$ 0.47$\dagger$ 0.11$\dagger$ 24.14$\dagger$ 46.3$\dagger$ 0.49$\dagger$ -0.44$\dagger$ 24.18$\dagger$
W4A8 Ours 11.60 0.59 0.33 26.08 16.7 0.55 0.11 25.37
SD-XL 768 FP16 - 0 1 0.49 26.51 0 1 0.69 29.74*/27.28*
W4A8 Q-Diffusion 44.07 - - 15.91 24.98 - - -9.81
W4A8 PTQ4DM 46.45 - - 15.92 28.28 - - -9.61
W4A8.4 PCR 18.27 0.48$\dagger$ -0.35$\dagger$ 23.85 18.25 0.52$\dagger$ -0.15$\dagger$ -4.03
W4A8.4 Ours 13.35 0.61 0.07 24.03 13.08 0.65 0.27 -3.08
SD3 512 FP16 - 0 1 0.88 26.25 0 1 0.91 25.86
W8A8 Ours 8.99 0.56 0.89 26.36 9.02 0.55 0.92 25.89
W4A16 NF4(g=128) 25.64 0.46 0.50 25.87 32.22 0.42 0.69 25.70
W4A16 NF4(g=64) 13.08 0.48 0.72 26.12 16.37 0.44 0.85 25.91
W4A8 Ours 9.77 0.52 0.90 26.35 11.03 0.47 0.95 26.09
1024 FP16 - 0 1 0.99 26.53 0 1 1.05 26.33
W8A8 Ours 8.53 0.67 0.96 26.54 9.07 0.64 1.04 26.07
W4A16 NF4(g=128) 32.12 0.56 0.61 25.78 33.61 0.52 0.79 25.89
W4A16 NF4(g=64) 12.97 0.58 0.86 26.23 15.03 0.55 0.97 26.14
W4A8 Ours 9.64 0.62 1.00 26.66 12.17 0.56 1.06 26.39
Figure 7. Stable Diffusion v1.4 512$\times$512 image generation using COCO prompts and Stable-Diffusion-Prompts.
Figure 8. Stable Diffusion v3 1024$\times$1024 image generation using COCO prompts and Stable-Diffusion-Prompts.
4.3. Objective Function
@@
We optimize quantized DDB $\epsilon_{q}$ to mimic the output of the floating-point DDB $\epsilon_{fp}$. Given the latent $\{x_{t1}\ldots x_{tn}\}$ at timestep $\{t_{1}\ldots t_{n}\}$, text embedding $\{p_{1}\ldots p_{n}\}$ from frozen text encoder, and precalculated projection $ep$, the output loss is defined as the mean squared error between the quantized and floating-point DDB outputs:
(9) $\mathcal{L}_{\text{out}}=\mathbb{E}\left[\left\|\epsilon_{fp}(x_{t_{1}\ldots t_{n}},p_{1\ldots n},e_{p})-\epsilon_{q}(x_{t_{1}\ldots t_{n}},p_{1\ldots n},e_{p})\right\|_{2}^{2}\right]$
where $\epsilon_{fp}$ and $\epsilon_{q}$ indicate the floating-point DDB and the quantized DDB, respectively.
@@
To perform inter-layer distillation, the loss function of the feature maps by the sensitive layers is added:
(10) $\mathcal{L}_{\text{sen}}=\mathbb{E}\left[\left\|f_{fp}^{s}(x_{t_{1}\ldots t_{n}},p_{1\ldots n},e_{p})-f_{q}^{s}(x_{t_{1}\ldots t_{n}},p_{1\ldots n},e_{p})\right\|_{2}^{2}\right]$
@@
5. Experiments
5.1. Experimental Setup
@@
Datasets. In this paper, we conduct experiments using two distinct datasets: COCO (Lin et al., 2014) and Stable-Diffusion-Prompts. We utilize prompts from the COCO2017 training dataset to construct the latent dataset. In terms of evaluation, the process is twofold. Following (Tang et al., 2024), firstly 5,000 prompts are selected from the COCO2017 validation dataset, which has been extensively employed in previous studies. Secondly, an additional 5,000 prompts from the Stable-Diffusion-Prompts dataset are used to assess the generalization capabilities of our quantized model in different prompt scenarios.
@@
Metrics. We evaluate the generative results of quantized models through three dimensions: (1) Similarity, (2) Quality, and (3) Text-image alignment. We compute the Fréchet Inception Distance between images generated by the quantized model and floating-point model (FID-to-FP (Tang et al., 2024)) to compute distributional similarity, alongside Structural Similarity (SSIM (Wang et al., 2004)) for pixel-level visual correspondence. Image quality is assessed using Image Reward (IR (Xu et al., 2023)), a human-preference-aligned metric that correlates with subjective ratings. Additionally, we use CLIP score (Hessel et al., 2021) to evaluate the matching degree between images and prompts. The evaluation code is adopted from  (Teng et al., 2023) and  (Chen and Mo, 2022).
@@
Baselines and implementation. We compare our proposed approach against advanced techniques: Q-diffusion (Li et al., 2023), PTQ4DM (Shang et al., 2023), PCR (Tang et al., 2024), and NF4 (Dettmers et al., 2023). Results are obtained from PCR (Tang et al., 2024) or reproduced. We employ the Stable Diffusion v1-4, Stable Diffusion v2-1, Stable Diffusion XL 1.0, and Stable Diffusion v3, both sourced from Hugging Face. We compare extensively with PCR (Tang et al., 2024) and align the quantization and generation settings with it. We use per-channel weight quantization and static per-tensor activation quantization. Except for special declaration, we adopt the default noise scheduler and Classifier-Free Guidance (CFG) scale. All experiments are conducted using a single NVIDIA A100.
5.2. Main Results
Quantitative results We demonstrate metric evaluation results in Tab. 2 across various models and precision levels.
on Stable Diffusion v1.4 & v2.1, Our method consistently outperforms PCR under the W4A8 configuration across all metrics, without employing the activation relaxation strategy in PCR (PCR relaxes 20% timesteps’ activation bitwidth to 10-bit, so it’s actually 8.4bit). Our approach achieves significantly higher similarity to FP16 baselines with negligible degradation in IR scores. Under W8A8, while our FID-FP is marginally higher than PCR’s W8A8.4, our method dominates other metrics.
For SDXL, we validate PCR’s observation that activation bitwidth relaxation is critical for text-image alignment (CLIP Score). Adopting PCR’s settings, our method still delivers substantial improvements over PCR, with empirical evidence showing that CLIP Score degradation correlates strongly with IR drop.
Additionally, we validate our method on MM-DiT based Stable Diffusion v3. Since previous methods haven’t tried on sd3, we establish NF4 (4-bit non-uniform weight quantization) as the baseline. Our W4A8 surpasses NF4’s W4A16 at both 512×512 and 1024×1024 resolutions, demonstrating scalability to advanced architectures.
Visual results. We show some corresponding qualitative comparisons on SD1.4 and SD3 in Fig. 7 and Fig. 8. Previous methods, when quantized to 4-bit, result in noticeable style changes in the generated images compared to those produced by the floating-point model. Such changes include but are not limited to, alterations in scene layout and facial features, loss of color and object, and the blending of multiple objects. In contrast, the images generated by our method are consistently of high fidelity.
5.3. Dataset Generation Analysis
@@
In this section, We will discuss the trade-off of some crucial hyperparameters in the latent dataset generation.
Table 3. Comparison of different sampling strategies for latent dataset generation.
@@
50steps/prompt 4000 6G 1.8h 9.44
1steps/prompt 20000 0.6G 4h 9.46
Figure 9. Comparison of Loss and FID-to-FP Curves under different number of prompts. The overfitting problem is obvious.
Table 4. Efficiency comparison on Stable Diffusion v1-4 and Stable Diffusion XL. (a) denotes sampling 50 steps per prompt, and (b) denotes sampling 1 step per prompt.
@@
Ours Gen Dataset (Once for all) (a) (b) (a) (b)
1.8h $\approx$4h 2.5h $\approx$5.5h
@@
More prompts or more timesteps? In the dataset generation process, we can randomly sample varying amounts of latent for each prompt. For comparison, two datasets have been constructed. The first dataset comprises 4000 prompts, with 50 latent per prompt. The second consists of 20000 prompts, each prompt with just 1 latent. As demonstrated in Tab. 3, despite the first dataset having 10$\times$ more latent, it exhibits a similar FID-to-FP. We can infer that a dataset with more prompts is more resistant to overfitting and has a smaller size but with a longer generation time. Given the comparable outcomes of the two strategies, we choose the latter dataset for smaller storage
Length of dataset. Given our limited training iterations, prompt quantity plays a critical role. Figure 9 clearly demonstrates the overfitting challenges with smaller datasets. To mitigate this, we employ default settings of 20,000 prompts for Stable Diffusion v1.4 and 10,000 prompts for Stable Diffusion XL, using 1-step sampling. As summarized in Table 4, our method achieves substantial training efficiency gains over PCR, particularly for larger architectures like Stable Diffusion XL.
5.4. Ablation Study
Tab. 5 presents our ablation study on Stable Diffusion v2-1 with W4A8 quantization. The ’Base’ configuration represents the original serial pipeline as illustrated in Fig. 1. We systematically evaluate each component’s contribution by gradually adding each of them. The results clearly show that each proposed component contributes meaningfully to the improvement of fidelity. Notably, the Serial-to-Parallel pipeline exhibits the most pronounced effect. Our method incorporates all these components effectively.
Table 5. Ablation results on the COCO prompts for Stable Diffusion v2-1 under W4A8 settings. ’S2P’ denotes the Serial-to-Parallel pipeline, ’Time’ denotes time information precalculation, ’Mstep’ denotes multi-step activation quantization, ’Distill’ denotes inter-layer distillation, and ’Freeze’ denotes selective freezing.